# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v1

@@ IV.4 In-depth Analysis of PQK Design Options
An overview of the model performance for each of these datasets across the different outer kernel functions per $O^{k}$, aggregated over the encoding circuits considered here, is provided in Appendix D.3, with Fig. A16 showing regression tasks and Fig. A17 illustrating classification problems. Here, we aim to quantify the influence of the respective choices on the model performance, while taking into account the corresponding impact of all other hyperparameters. For this purpose, we aggregate all simulation results per dataset and conduct (semi-) partial correlation analyses (cf. Appendix. A for details). The results are summarized in Tab. 1. Note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
@@
To grasp these results illustratively, we once more refer to the supporting Figs. A16 and A17 and the corresponding explanations there. From the pure numbers, we directly realize that both $\kappa$ and the choice of $O^{k}$ for calculating the respective projected quantum circuits exhibit statistically significant correlations. For the outer kernel, no universal statement can be made a priori regarding which one is generally suitable for effective QKM design—at least not within the scope of our study. The only exception is that for the classification tasks considered here, the RationalQuadratic kernel is not a good choice. Furthermore, it is worth noting that the commonly used vanilla PQK in the literature, i.e., with a Gaussian kernel, is not automatically the best choice.
In relation to a proper choice of the measurement operator $O^{k}$, again there appears to be no ad-hoc answer towards which one is best suited to achieve good model performance. Only $X_{k=1}$ seems to be generally insufficient, wherefore we frequently observed training kernel Gram matrices filled entirely with ones, thus rendering them impractical for learning. In contrast, Tab. 1 suggests a potential improvement in model performance when utilizing the $P^{1+2}$ measurement operator; cf. Eq. (18).
Apart from that, the results indicate that the definition of the projected quantum circuits via $O^{k}$, as well as the choice of $\kappa$ should be considered as additional hyperparameters for PQK model design. Furthermore, the varying results between each dataset may suggest that this hyperparameter search and thus the fundamental definition of PQKs should be again dataset-driven; analogous to the data encoding circuit architecture search. A systematic investigation of this is however still lacking. But in case this is being confirmed, the correlating significance in relation to the promised benefits of PQKs in regard to circumventing the curse of dimensionality of FQKs [16] requires further exploration.
In Ref. [34] the authors have raised the question whether the projected quantum circuit is not so much responsible for learning, but rather the subsequent outer (Gaussian) kernel applied to the features computed by Eq. (II). To this end, referring first to Tab. 1, it becomes evident, as previously discussed, that both the choice of $O^{k}$ (and thus the projected quantum circuit itself), and the outer kernel $\kappa$ are crucial for achieving sufficiently good model performance, depending on the dataset. Therefore, the question of which factor is more important cannot be definitively answered based on this analysis alone. Consequently, we further list all (statistically significant) correlations of the various hyperparameters of the QKMs for a given outer kernel, both among themselves and with model performance in Appendix D.3 in Tab. 2. Thereof, it again becomes clear that depending on the dataset and outer kernel function, both hyperparameters that contribute to the definition of the projected quantum circuits (i.e., $n_{\mathrm{layers}}$ and $w_{\mathrm{e}}$) and those solely associated with the outer kernel exhibit statistically significant correlations.
@@
Thus, calculating $d(G|G^{\prime})$ also defines a difference measure between projected quantum circuit contribution $F$ and $F^{\prime}$, which moreover also holds for the Matérn and RationalQuadratic outer kernels, cf. Eqs. (12) and (13), respectively. Figure 12 exemplifies this for the QFMNIST ($n=8$) dataset for Gaussian outer kernel function.
@@
Figure 12: Comprehensive insight into various mechanisms of PQKs exemplified for the QFMNIST dataset with $n=8$ principal components and Gaussian outer kernel function. The central heatmap plot illustrates the $d(G|G^{\prime})$ distance measure according to Eq. (19), where each section correpsonds to a family of projected quantum circuits as defined by a given measurement operator $O^{k}$. The sections are further subdivided into the different encoding circuits in the order as the appear in Appendix B. The plot at the top details $\mathrm{Var}(F)$; cf. Eq. (20). The close-up images on the very right correspond to examples of $F$ and $F^{\prime}$ matrices that lead to $d(G|G^{\prime})\to 0$ and $d(G|G^{\prime})\neq 0$, respectively. The $F$ matrices can be viewed as illustrations of embeddings into the corresponding Hilbert space. The histogram of MSEs right to the central heatmap helps to identify Gram matrices that lead to good/bad performances, respectively. The $w_{\mathrm{e}}$- and $\gamma$-plots on the left and on the bottom, respectively helps to understand the machanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large.
@@
Each section in the central $d(G|G^{\prime})$ heat map plot corresponds to a family of projected quantum circuits defined by a given measurement operator $O^{k}$ and is further subdivided into the different encoding circuits in the order as they appear in Appendix B. There are both regions with finite distances between two Gram matrices and those with vanishing differences. This can be explained considering the $\mathrm{Var}(F)$ plot at the top as well as the enlarged examples of the $F$ matrices leading to the respective $d(G|G^{\prime})$ entries. Vanishing $d(G|G^{\prime})$ distances correspond to Gram- and $F$-matrices that show favourable variance behavior or, in other words, lead to similar embedding $F$ and $F^{\prime}$ in the corresponding Hilbert space (upper zoom-in plot). In contrast, finite distances (lower zoom-in plot) result from one Gram-/$F$-matrix being ill-conditioned, i.e., $\mathrm{var}(G)\to 0$ and thus $\mathrm{var}(F)\to 0$ or vice versa. Consequently, these matrices correspond to weakly performing (in the sense of MSE scores for regression) models, while the others show better MSE scores, as can be inferred from the respective plot on the right. By considering the $w_{\mathrm{e}}$- and $\gamma$-plots on the left and on the bottom, respectively helps to understand the machanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large. With this mechanism we can a posteriori explain the origin of bad performing PQK models in Figs. 6 and 7.
The analogous plots for all datasets as well as for the remaining outer kernel functions are detailed in the Appendix D in Figs. A18 to A31.
@@
Eventually, this analysis (as well as Eq. (20) and analogous relations for Matérn and RationalQuadratic outer kernels) confirms our hypothesis that there is a synergy and dependence between projected quantum circuits and outer kernel as well as QKM hyperparameters. As such, also given the insights from Tabs. 1 and 2, we conclude that both parts are responsible for learning.
@@ V Discussion
In this benchmarking study, we systematically analyzed various design criteria of QKMs based on both FQKs and PQKs for several regression and classification tasks. In order to gain deeper insights into the functional principles underlying QKMs and derive universal insights, we conducted hyperparameter optimizations and training of more than $20,000$ models. However, although we attempted to encompass multiple different dataset families, the effective number of distinct families, the size of each dataset, and the specifics of splitting them into training and test sets represents a restriction, which limits the generalizability of our results.
@@
The study of hyperparameter importances for effective model tuning illustrated that regularization, length scale (for PQKs) and data prescaling hyperparameters, are mostly driving the successful training of QKMs within our simulations. These results further indicated that the number of qubits and layers of the respective data encoding circuits are less important, both for FQK and PQK approaches. Given that for increasing dataset complexity, in particular for the two classification tasks considered here, the number of qubits and layers for successful QKM training may become more relevant, the pronounced importance of these classical hyperparameters found here can probably be attributed to the (low) difficulty of some of the dataset configurations. In this respect, it is known that quantum models can be written as partial Fourier series in the data, where the accessible frequencies are eventually determined by the number of qubits and layers of the underlying data encoding circuits [14]. Moreover, while Ref. [11] outlines the Fourier representation of quantum kernels, it remains an interesting question how this can be leveraged for the QKM design as well as for the identification of promising dataset properties.
@@
Moreover, our analysis of model performances demonstrated that the average feature correlation to the targets can serve as a useful difficulty measure for predicting test performance scores. We generally found that as dataset complexity increases, the higher/lower the MSE/ROC-AUC metrics, respectively. Interestingly, for regression tasks the particular choice of kernel method, i.e., QKRR and QSVR approaches is not really important as, on average, they are performing equally well. Similarly, our experiments across both regression and classification tasks, show no significant performance differences between FQK and PQK, even though we simulate up to 15 qubits. In this range, the first effects of the limited trainability of FQKs due to the exponential concentration phenomenon [28] should appear, and one might expect evidence of better PQK performance. The absence of this observation however, may be explained by bandwidth tuning (here, we optimize $w_{\mathrm{e}}$), whereby generalizability can be established, but, as shown in Refs. [29, 37], at the expense of becoming classically tractable. To properly understand the connections between bandwidth tuning, exponential concentration and classical tractability in terms of the geometric difference metric [16] will be part of a future work. Our Spearman correlation analyses further validated the importance of classical hyperparameters, aligning with our findings from the hyperparameter importance study.
@@
The survey of different data encoding strategies in terms of embedding some or all features redundantly, revealed that for regression tasks based on FQKs it is beneficial to redundantly encode all features on a number of qubits that is given by integer multiples of the number of features. For PQKs redundant embedding appears to be negligible. For the classification tasks considered in this work, we observed best performances for both FQK and PQK for a number of qubits corresponding to non-integer multiples of the number of features.
Our exploration of PQK design options, including the effect of various measurement operators and outer kernel functions, hinted at that projected quantum circuits and classical hyperparameters, particularly regularization strengths, data prescaling, and length scale parameters, work synergistically to drive learning. However, these findings as well as those resulting from detailed correlation analyses indicate that our understanding of “quantumness” of QKMs can be challenged even further, apart from the already mentioned entanglement problem. At this we note that investigating shadow kernels or linear PQKs [16] accordingly remains open.
Concerning proper datasets for QKM tasks, our results for QFMNIST, which can be seen as artificially engineered quantum data, may be interpreted such that there is some weak indicative that the actual quantum part of QKMs, i.e., the number of qubits and layers of the data encoding circuits as well as the choice of the measurement operator in case of PQK become slightly more important.
In conclusion we believe that this study, substantially contribute to better understand what distinguished good from bad QKMs. For the datasets we investigated in this work, we showed that in most cases it remains unresolved to clearly reveal the quantum contribution and its importance for learning. At most for PQKs we could reveal that classical and quantum parts can contribute likewise. In our view, these findings corroborate the need to tackle QML research from two sides, i.e.  identify requirements for datasets with potential quantum advantage and explore how corresponding model design should look like.
@@ Appendix D Detailed Results
Figure A10: Analogous illustration to Fig. 6 with associated training performance scores. Results for each QKM and dataset are aggregated across data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$, yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays corresponding findings for regression datasets. Here, (a) shows the Friedman dataset family with datsets for #$\#\mathrm{features}n=[5,15]$. In (b) we give results corresponding to the QFMNIST dataset family with datasets having number of principal components $n=[2,15]$. The lower panel shows the classification tasks. The results in (c) correspond to the two curves diff family with $d=4$ and degree $D=[2,20]$. In (d) we present the hidden manifold diff family with $d=4$ and manifold dimension $m=[2,20]$.
D.1 Details on model performance and correlation analysis
In Sec. IV.2 we mention that both regression and classification datasets exhibit a propensity for overfitting. The Friedman and QFMNIST regression datasets in Figs. A10 (a) and (b), respectively, reveal that each dataset (in terms of $n$) shows lower whiskers with $\mathrm{MSE}\to 0$. Comparing to the corresponding test performance scores in Figs. 6 (a) and (b), this may also indicate signatures of overfitting especially for datasets with larger $n$. In this respect, one should carefully investigate different training- and test set sizes, which represents a limiting factor of this study. The corresponding training performance scores for the two curves diff and hidden manifold diff datasets are displayed in Figs. A10 (c) and (d), respectively.
When investigating the impact of data encoding circuits on QKM model performance, we merely presented results of the Friedman (regression) and the two curves diff (classification) datasets in Fig. 7. Here, we provide analogue findings in Fig. A11 for the QFMNIST and the hidden manifold diff datasets.
Figure A11: Complementing illustration to Fig. 7 of the main text with the datasets not shown therein. Upper panel: Regression QFMNIST. With (a) Test performance aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel for the dataset with number of features $n=15$ and (b) test performance scores aggregated over all datasets in the QFMNIST family and over respective $n_{\mathrm{layers}}$. Lower panel: hidden manifold diff. With (c) Test performance aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with manifold dimension $m=13$ and (d) test performance scores aggregated over all datasets in the hidden manifold diff family and over respective $n_{\mathrm{layers}}$.
Finally, we summarize those results of the correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameter themselves, that are not shown in the main text in Sec. IV.2 in the following Fig. A12.
@@ D.1 Details on model performance and correlation analysis
D.2 Details on data encoding strategies
The results for testing the two different encoding strategies for the Friedman dataset for $n=5$ features as outlined in the main text, cf. Sec. IV.3 and Fig. 9, are displayed in Fig. A13. Contrary to the QFMNIST results in Fig. 10, in all cases, except for FQK, option 2, most accurate results are obtained for $n_{\mathrm{qubits}}=12$. Here, FQK and PQK approaches show similar behavior. The second encoding option for FQK appears to be less sensitive for the respective feature redundancy. Notably, that the redundancy does not have to be an integer multiple of the original number of features.
Figure A13: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the Friedman regression dataset $n=5$ features. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.

### v2

@@ IV.5 Analysis of PQK Design Options
We aim to quantify the influence of the respective choices on the model performance, while taking into account the corresponding impact of all other hyperparameters. For this purpose, we aggregate all simulation results per dataset and conduct (semi-) partial correlation analyses (cf. Appendix. A for details). The results are summarized in Tab. 1. Note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, the opposite is the case for correlations with ROC-AUC scores. An overview of the model performances used for calculating the values in the table is provided in Appendix D.5 in Figs. A16 and A17.
@@
From the table we directly realize that both $\kappa$ and the choice of $O^{k}$ for defining the respective projected quantum circuits exhibit statistically significant correlations. For the outer kernel, no universal statement can be made a priori regarding which one is generally suitable for effective QKM design—at least not within the scope of our study. The only exception is that for the classification tasks considered here, the RationalQuadratic kernel is not a good choice. Furthermore, it is worth noting that the commonly used vanilla PQK in the literature, i.e., with a Gaussian kernel, is not automatically the best choice.
Regarding a proper choice of the measurement operator $O^{k}$, Tab. 1 suggests $X_{k=1}$ to be insufficient for all regression and classification tasks considered here. In this case, we frequently observed training kernel Gram matrices entirely filled with ones, thus rendering them impractical for learning. In contrast, utilizing the $P^{1+2}$ measurement operator, cf. Eq. (18), points towards an overall improvement in model performance (within the scope of this study).
Apart from that, the results indicate that the definition of the projected quantum circuits via $O^{k}$, as well as the choice of $\kappa$ should be considered as additional hyperparameters for PQK model design. A systematic investigation of this is however still lacking.
In Ref. [34] the authors have raised the question whether the projected quantum circuit is not so much responsible for learning, but rather the subsequent outer (Gaussian) kernel applied to the features computed by Eq. (II). Referring to Tab. 1 per dataset, it becomes evident that both the choice of $O^{k}$ (and thus the projected quantum circuit itself), and the outer kernel $\kappa$ are crucial for achieving sufficiently good model performance. Therefore, the question of which factor is more important cannot be definitively answered based on this analysis alone. Consequently, we further list all (statistically significant) correlations of the various hyperparameters of the PQK QKMs for a given outer kernel, both among themselves and with model performance in Appendix D.5 in Tab. 2. Thereof, it again becomes clear that depending on the dataset and outer kernel function, both hyperparameters that contribute to the definition of the projected quantum circuits (i.e., $n_{\mathrm{layers}}$ and $w_{\mathrm{e}}$) and those solely associated with the outer kernel exhibit statistically significant correlations.
@@
Thus, calculating $d(G|G^{\prime})$ also defines a difference measure between projected quantum circuit contributions $F$ and $F^{\prime}$, which moreover also holds for the Matérn and RationalQuadratic outer kernels, cf. Eqs. (12) and (13), respectively. Figure 12 exemplifies this for the QFMNIST ($d=8$) dataset for Gaussian outer kernel function.
@@
Figure 12: Comprehensive insight into various mechanisms of PQKs exemplified for the QFMNIST dataset with $d=8$ principal components and Gaussian outer kernel function. The central heatmap plot illustrates the $d(G|G^{\prime})$ distance measure according to Eq. (19), where each section correpsonds to a family of projected quantum circuits as defined by a given measurement operator $O^{k}$. The sections are further subdivided into the different encoding circuits in the order as they appear in Appendix B. The plot at the top details $\mathrm{Var}(F)$; cf. Eq. (20). The close-up images on the very right correspond to examples of $F$ and $F^{\prime}$ that lead to $d(G|G^{\prime})\to 0$ and $d(G|G^{\prime})\neq 0$, respectively. The heatmap visualization of $F$ can be viewed as illustrations of embeddings into the corresponding Hilbert space. The histogram of MSEs right to the central heatmap helps to identify Gram matrices that lead to good/bad performances (with bad defined as MSE $>0.1$, indicated by orange dashed line), respectively. The $w_{\mathrm{e}}$- and $\gamma$-plots on the left and on the bottom, respectively helps to understand the machanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large.
@@
Each section in the central $d(G|G^{\prime})$ heatmap plot corresponds to a family of projected quantum circuits defined by a given measurement operator $O^{k}$ and is further subdivided into the different encoding circuits in the order as they appear in Appendix B. There are regions with finite distances between two Gram matrices and regions with vanishing differences. This can be explained considering the $\mathrm{Var}(F)$ plot at the top as well as the zoomed-in examples of the $F$ heatmaps on the very right. Vanishing $d(G|G^{\prime})$ distances correspond to Gram- and $F$-matrices that show favourable variance behavior or, in other words, lead to similar embeddings $F$ and $F^{\prime}$ in the corresponding Hilbert space (upper zoom-in plot). In contrast, finite distances (lower zoom-in plot) result from one Gram-/$F$-matrix being ill-conditioned, i.e., $\mathrm{var}(G)\to 0$ and thus $\mathrm{var}(F)\to 0$ or vice versa. Consequently, these matrices correspond to weakly performing models, while the others show better MSE scores as can be observed from the respective histogram plot on the right. Considering the $w_{\mathrm{e}}$- and $\gamma$-histograms on the left and on the bottom helps to understand the mechanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large. With this mechanism we can a posteriori explain the origin of bad performing PQK models in Figs. 4 and 8. The analogous plots for the QFMNIST($d=8$) dataset for the remaining outer kernel functions are detailed in the Appendix D.
@@
This analysis confirms our hypothesis that there is a synergy and interdependence between projected quantum circuits and outer kernels as well as QKM hyperparameters. As such, we conclude that PQK-based QKMs require careful examination of outer kernels and suitable measurement operators to perform well on a given dataset.
@@ V Discussion
In this benchmarking study, we systematically analyzed various design criteria of QKMs that are based on FQKs and PQKs for several regression and classification tasks. In order to derive universal insights, we conducted hyperparameter optimizations and training of more than $20,000$ models.
@@
The study of hyperparameter importances for effective model tuning illustrates that regularization, outer kernel length scale (for PQKs) and data prescaling hyperparameters, are mostly driving the successful training of QKMs within our simulations. These results further indicate that the number of qubits and layers of the respective data encoding circuits are less important, both for FQK and PQK approaches. However, for increasing dataset complexity, in particular for the two classification tasks considered here, our findings indicate an increased importance of the number of qubits and layers. Hence, identifying and applying sufficiently complex datasets may pave the way towards a deeper understanding of quantum kernel methods. In this respect we note that it is known that quantum models can be written as partial Fourier series in the data, where the accessible frequencies are eventually determined by the number of qubits and layers of the underlying data encoding circuits [14]. Moreover, Ref. [11] outlines the Fourier representation of quantum kernels. However, it remains an interesting open question how this can be leveraged for the QKM design as well as for the identification of promising dataset properties.
@@
Furthermore, our analysis of the model performances demonstrates that the average feature correlation to the targets can serve as a useful difficulty measure for predicting test performance scores. Interestingly, for regression tasks the particular choice of kernel method, i.e., QKRR and QSVR approaches is not really important as, on average, they are performing equally well. Similarly, our experiments across both regression and classification tasks show no significant performance differences between FQK and PQK, even though we simulate up to 15 qubits. In this range, the first effects of the limited trainability of FQKs due to the exponential concentration phenomenon [28] should appear, and one might expect evidence of better PQK performance. The absence of this observation however, may be explained by bandwidth tuning (here, we optimize $w_{\mathrm{e}}$), whereby generalizability can be established, but, as shown in Refs. [29, 37], at the expense of becoming classically tractable. To properly understand the connections between bandwidth tuning, exponential concentration and classical tractability in terms of the geometric difference metric [16] will be part of a future work. Our Spearman correlation analyses further validates the importance of classical hyperparameters, aligning with our findings from the hyperparameter importance study.
@@
The survey of different data encoding strategies in terms of embedding some or all features redundantly reveals that for tasks involving datasets with lower complexity, based on FQKs, it is beneficial to redundantly encode all features on a number of qubits that is given by integer multiples of the number of features. For PQKs, redundant embedding appears to be negligible. For tasks involving datasets with higher complexity, we observed the best performances for both FQK and PQK for a number of qubits corresponding to non-integer multiples of the number of features.
Our exploration of PQK design options reveals that projected quantum circuits, outer kernel length scale parameters, regularization strengths, and data prescaling hyperparameters drive the learning capability of PQK-based QKMs synergistically. At this, we note that investigating shadow kernels or linear PQKs [16] accordingly remains open.
Beyond that, we again emphasize the need for proper and sufficiently complex datasets to obtain deeper insights into quantum kernel models. In regard to PQKs, this may promote our understanding of the impact of the choice of the measurement operator and the outer kernel. Related to this, it is worth mentioning again that although we encompass multiple different dataset families, the effective number of distinct families, the size of each dataset, and the specifics of splitting into training and test sets represents a restriction, which limits the generalizability of our results.
In conclusion we believe that this study substantially contribute to better understand what distinguishes good from bad QKMs. For the datasets we investigated in this work, we showed that in most cases it remains unresolved to clearly reveal the quantum contribution and its importance for learning. At most for PQKs we could reveal that classical and quantum parts can contribute likewise. In our view, these findings corroborate the need to tackle QML research from two sides, i.e.  identify requirements for datasets with potential quantum advantage and explore how corresponding model design should look like.
@@ Appendix D Detailed Results
Figure A10: Analogous illustration to Fig. 4 with associated training performance scores. Results for each QKM and dataset are aggregated across data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$, yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays corresponding findings for regression datasets. Here, (a) shows the Friedman dataset family with datsets for #$\#\mathrm{features}d=[5,15]$. In (b) we give results corresponding to the QFMNIST dataset family with datasets having number of principal components $d=[2,15]$. The lower panel shows the classification tasks. The results in (c) correspond to the two curves diff family with $d=4$ and degree $D=[2,20]$. In (d) we present the hidden manifold diff family with $d=4$ and manifold dimension $m=[2,20]$.
D.1 Details on Model Performance
In Sec. IV.1 we mention in connection with outliers in Fig. 4 that both regression and classification datasets exhibit a propensity for overfitting. The Friedman and QFMNIST regression datasets in Figs. A10 (a) and (b), respectively, reveal that each dataset shows lower whiskers with $\mathrm{MSE}\to 0$. Comparing to the corresponding test performance scores in Figs. 4 (a) and (b), this may also indicate signatures of overfitting especially for datasets with larger $d$. The corresponding training performance scores for the two curves diff and hidden manifold diff datasets are displayed in Figs. A10 (c) and (d), respectively. As discussed in the main text, since we investigate a large variety of combinations, it is unavoidable that some of them result in models that are too expressive for some tasks. Moreover, we point out that one should carefully investigate different training- and test set sizes in future studies.
D.2 Additions to Hyperparameter Correlation Analysis
To complement the report of correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves in Sec. IV.2, we provide a summary of all other dataset families not shown in the main text in Fig. A11.
@@ D.2 Additions to Hyperparameter Correlation Analysis
D.3 Details on the Influence of Encoding Circuits
When investigating the impact of data encoding circuits on QKM model performance in Sec. IV.3, we merely presented results of the Friedman (regression) and the two curves diff (classification) datasets in Fig. 8. Here, we provide analogue findings in Fig. A12 for the QFMNIST and the hidden manifold diff datasets.
Figure A12: Complementing illustration to Fig. 8 of the main text with the datasets not shown therein. Upper panel: Regression QFMNIST. With (a) Test performance aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel for the dataset with number of features $d=15$ and (b) test performance scores aggregated over all datasets in the QFMNIST family and over respective $n_{\mathrm{layers}}$. Lower panel: Classification hidden manifold diff. With (c) Test performance aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with manifold dimension $m=13$ and (d) test performance scores aggregated over all datasets in the hidden manifold diff family and over respective $n_{\mathrm{layers}}$.
D.4 Details on the Influence of Data Encoding Strategies
The results for testing the two different encoding strategies for the Friedman dataset for $d=5$ features as outlined in the main text, cf. Sec. IV.4 and Fig. 9, are displayed in Fig. A13. Contrary to the QFMNIST results in Fig. 10, in all cases, except for FQK, option 2, most accurate results are obtained for $n_{\mathrm{qubits}}=12$. Here, FQK and PQK approaches show similar behavior. The second encoding option for FQK appears to be less sensitive for the respective feature redundancy. Notably, that the redundancy does not have to be an integer multiple of the original number of features.
Figure A13: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the Friedman regression dataset $d=5$ features. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.